# Machine-Learning Estimation of Energy Fractions in MHD Turbulence Modes

### v1

@@ 5.1 Comparison with earlier work
Furthermore, recent studies have shown that 21,cm H I thin-channel maps, when combined with the Galactic rotation curve, can be used to reconstruct the 3D distribution of the Galactic magnetic field (Hu & Lazarian, 2023; Schmaltz et al., 2024; Hu, 2025c). A similar idea could be incorporated into our ML framework, potentially enabling the reconstruction of the 3D distribution of turbulence-mode energy fractions throughout the ISM.
5.2 Implications for studying cosmic ray transport and star formation
@@
6 Conclusion
In this study, we investigated the kinetic-energy fractions of the three fundamental MHD turbulence modes, Alfvén, slow, and fast modes, under both isothermal and multiphase ISM conditions. We further developed a method that employs a conditional Residual Neural Network (ResNet) together with three distinct spectroscopic feature maps to estimate these mode fractions. This approach leverages the morphological differences imposed by distinct MHD modes and presents a novel method for determining mode composition from spectroscopic observations. Our main findings are as follows:
@@ 6 Conclusion
    Training the model on both isothermal and multiphase simulations shows that fast-mode predictions using channel maps achieve the lowest errors. For seen data, the mean prediction error is approximately 0.05, while for unseen data it increases to $\sim$0.1, indicating consistent performance across different thermal conditions.

### v2

@@ V.1 Comparison with earlier work
Furthermore, recent studies have shown that 21 cm H I thin-channel maps, when combined with the Galactic rotation curve, can be used to reconstruct the 3D distribution of the Galactic magnetic field (Hu & Lazarian, 2023b; Schmaltz et al., 2024; Hu, 2025b). A similar idea could be incorporated into our ML framework, potentially enabling the reconstruction of the 3D distribution of turbulence-mode energy fractions throughout the ISM.
V.2 Implications for studying cosmic ray transport and star formation
@@
V.3 Limitations
Several limitations merit consideration. First, both the training and validation datasets are derived from numerical simulations and are therefore subject to numerical artifacts. In particular, simulations inevitably introduce numerical dissipation of turbulence, whereas such dissipation is not present in real observations. Second, observational data are affected by instrumental noise and finite beam size, neither of which is included in the current simulations. We do not expect noise to significantly degrade the model performance: in our earlier studies on predicting the sonic Mach number and magnetic field strength (Schmaltz et al., 2025; Hu, 2025b), the machine-learning framework remained robust for signal-to-noise ratios exceeding three.
In addition, the experiments presented here explore only a restricted region of parameter space. All simulations are performed in the strong-field, sub-Alfvénic ($M_{A}\leq 1$) regime, whereas the interstellar medium may be super-Alfvénic ($M_{A}>1$) on large scales. Although the ResNet exhibits robust performance across both isothermal and multiphase conditions, its ability to generalize to super-Alfvénic turbulence has yet to be assessed.
VI Conclusion
In this study, we investigated the kinetic-energy fractions of the three fundamental MHD turbulence modes, Alfvén, slow, and fast modes, under both isothermal and multiphase ISM conditions. We further developed a method that employs a conditional Residual Neural Network (ResNet) together with three distinct spectroscopic maps to estimate these mode fractions. This approach leverages the morphological differences imposed by distinct MHD modes and presents a novel method for determining mode composition from spectroscopic observations. Our main findings are as follows:
@@ VI Conclusion
    Training the model on both isothermal and multiphase simulations shows that fast-mode predictions using channel maps achieve the lowest errors. The mean relative normalized errors of approximately 0 and the standard deviation of 0.01 - 0.02 for seen data and 0.1 - 1.8 for unseen data.
@@
Appendix A Compressive multiphase simulation
[…]
We further evaluated the ResNet performance in estimating mode energy fractions using this same compressible simulation (Fig. 8). The predicted mode fractions recover the ground-truth values. This behavior persists across a broader set of input channel maps, indicating that the ResNet architecture generalizes well to both solenoidally driven and compressively driven MHD turbulence when estimating mode energy fractions.